# Roundtable Policy: Improving Scientific Reasoning and Narratives through Confidence-Weighted Consensus of LLMs

### v1

@@ 5 Related Work
LLMs for Scientific Discovery. Large language models (LLMs) have rapidly expanded the scope of data-driven science, from AlphaFold’s breakthroughs in structural biology (Jumper et al., 2021) to recent applications across mathematics (Yue et al., 2024), chemistry (Yu et al., 2024), biology (Labrak et al., 2024), and geography (Li et al., 2023). Their strength lies in structured reasoning and seamless integration with external tools, enabling automated workflows (Cao, 2017). Yet, complex multi-step reasoning often requires specialized prompting, tool-augmented inference, or fine-tuning, motivating the rise of agentic systems that combine reasoning with external solvers and databases (Boiko et al., 2023; Lu et al., 2024).
Reasoning and Factuality in Language Models. Improving reasoning and factuality in LLMs has been studied extensively. Reasoning can be enhanced through prompting methods such as scratchpads (Nye & et al., 2021), chain-of-thought demonstrations (Wei et al., 2022; Kojima et al., 2022; Reynolds & McDonell, 2021), verification and self-consistency (Cobbe et al., 2021; Wang et al., 2022), and reflection (Shinn et al., 2023; Madaan & et al., 2023), as well as fine-tuning approaches (Rajani et al., 2019; Zelikman et al., 2022; Lewkowycz et al., 2022). Factuality, in turn, has been strengthened through RLHF (Christiano et al., 2017; Ziegler & et al., 2019; Ouyang et al., 2022), dataset curation (Lee & et al., 2022), retrieval-based augmentation (Guu et al., 2020), and likelihood-based calibration (Kadavath & et al., 2022).
Multi-Agent and Ensemble Approaches. Beyond single-model prompting, researchers have explored ensemble and multi-agent strategies to enhance reliability. Early work shows that self-consistency (Wang et al., 2022) and majority voting (Li et al., 2022; Cobbe et al., 2021; Thoppilan et al., 2022) can mitigate errors by aggregating diverse reasoning chains. Building on this, multi-agent debate frameworks (Irving et al., 2018; Du & et al., 2023; Madaan & et al., 2023) and collaborative dialogue methods (Chan & et al., 2023; Wang & et al., 2023) let models iteratively critique and refine one another’s outputs. Related lines of work also investigate compositional generation, combining outputs from multiple models (Du et al., 2020; Liu & et al., 2022; Alayrac & et al., 2022; Zeng & et al., 2022). While these methods demonstrate the value of collective intelligence, they often rely on flat aggregation or emergent convergence. In contrast, our approach introduces a structured consensus mechanism that explicitly weights models by demonstrated reliability, producing interpretable and auditable signals of trust.
Benchmarks for LLM-Based Agents. Evaluating agentic LLMs has inspired benchmarks ranging from synthetic tool-use tasks (e.g., TaskBench (Shen et al., 2024)) to domain-specific challenges such as SWE-Bench (Jimenez et al., 2024), BioCoder-Py (Tang et al., 2024), and MLAgentBench (Huang et al., 2024). More recently, datasets sourced from scientific literature (Majumder et al., 2024) aim to capture authentic workflows. Our work complements these efforts by introducing a domain-authentic benchmark in materials science that stresses both mechanistic understanding and property prediction, thereby testing agents under both interpretative and predictive dimensions.
Machine Learning and Agents for Materials Science. Machine learning has already transformed materials science by enabling scalable property prediction, compound discovery, and structure–property analysis (Jain et al., 2013; Butler et al., 2018), with advances in reinforcement learning for inverse design (Segler et al., 2018), generative modeling (Granda et al., 2018), and multi-modal integration of simulations and experiments (Pyzer-Knapp et al., 2022). These developments reduce experimental costs and provide interpretable insights, but LLM-based agents remain relatively underexplored in this domain. The inherent complexity of materials science—multiscale interactions, vast compositional spaces, and intricate synthesis pipelines—calls for agent frameworks that unify predictive modeling with experimental feedback. Early systems in autonomous discovery (Gregoire et al., 2023; Maurizi et al., 2024) show promise, but important gaps remain in linking reasoning, planning, and uncertainty quantification. Our work positions materials science as a natural testbed for stress-testing multi-agent reasoning frameworks such as Roundtable Policy, bridging structured evaluation and open-ended scientific narrative.
6 Conclusion & Future Directions
We presented the Roundtable Policy, a inference-time reasoning framework that fuses outputs from multiple LLMs through confidence-weighted aggregation without finetuning base models. Unlike single-model prompting or naive ensemble methods such as majority voting, our approach does not treat all models equally nor rely on opaque emergent convergence, as in debate-style methods. Instead, it introduces a principled mechanism that explicitly encodes historical reliability and uncertainty, producing consensus answers that are not only more accurate but also transparent and auditable. Across structured MultiTask benchmarks and open-ended SingleTask proposal writing, we showed that Roundtable Policy consistently improves factual reliability, reasoning consistency, and narrative coherence, highlighting its value as a trustworthy tool for scientific reasoning.
Looking forward, several avenues can extend this framework. First, adaptive arbitrator strategies could allow the confidence-weight table to evolve dynamically with task complexity, reallocating trust in real time. Second, coupling the fusion process with reinforcement learning and human or experimental feedback may enable autonomous optimization of weight assignments through trial and error. Third, incorporating richer feedback signals—such as calibrated uncertainty, tool-augmented checks, or domain-specific rubrics—could further strengthen robustness in complex scientific workflows.
More broadly, the principles underlying Roundtable Policy generalize beyond the settings studied here. By tailoring consensus mechanisms to domain-specific evaluation criteria, the framework can support applications ranging from drug discovery and computational chemistry to astrophysics and other data-intensive sciences. In this sense, Roundtable Policy is not merely another ensemble method, but a scalable and interpretable paradigm for collective intelligence in LLMs—one that advances the rigor, reliability, and accountability of AI-assisted scientific discovery.
Ethics Statement
Our method significantly enhances the performance of sequential model editing, making it invaluable for updating and managing knowledge in real-world applications. While the ability to directly modify stored knowledge introduces potential risks, such as the introduction of false or harmful information, we strongly urge researchers to implement strict validation and oversight to ensure the ethical use of these techniques. Nevertheless, the original goal of model editing is positive, aiming to facilitate efficient updates of large models in the future. Therefore, we encourage researchers to leverage this technology responsibly and with care.
Reproducibility
To ensure the reproducibility of our findings, detailed implementation instructions for Roundtable Policy can be found in Appendix. Additionally, the source code is publicly available at https://ai4scidisc.github.io/roundtable/. These measures are intended to facilitate the verification and replication of our results by other researchers in the field.
Acknowledement
We gratefully acknowledge the MIT SuperCloud and Texas Advanced Computing Center (TACC) at The University of Texas at Austin for providing computational resources that have contributed to the research results reported within this paper. JL acknowledges support by the Defense Advanced Research Projects Agency (DARPA) under Agreement No. HR00112490369. The content is solely the responsibility of the authors and does not necessarily represent the official views of the funding entities.
@@ A.1 Symbols and Abbreviations
: the number of player agents (pretrained LLMs generating candidate responses). In our experiments .
@@
: the fusion agent, which synthesizes a final consensus answer from all player responses using the confidence-weight table.
@@
: the corresponding reference answers (available only in MultiTask).
@@
: the fused consensus response in round .
| Setting | Player ID | Abbreviation | Full Name | Citation |
MultiTask |
001 | GPT-4o | GPT-4o | OpenAI (2024b) |
| 002 | Grok-2 | Grok-2 | xAI (2024) | |
| 003 | Claude-3.5H | Claude-3.5 Haiku | Anthropic (2024) | |
| 004 | GPT-4.5p | GPT-4.5-preview | OpenAI (2024a) | |
| 005 | DS-chat | DeepSeek-Chat | AI (2024) | |
| 006 | Claude-3.7S | Claude-3.7 Sonnet | Anthropic (2025) | |
| 007 | Llama-3.3 | Llama-3.3 | AI (2025) | |
| 008 | Qwen+ | Qwen-plus | Team (2024) | |
| 009 | Gemini-2.0 | Gemini-2.0 | DeepMind (2024) | |
SingleTask |
001 | GPT-4o | GPT-4o | OpenAI (2024b) |
| 002 | GPT-o3 | GPT-o3 | OpenAI (2024d) | |
| 003 | GPT-4.1 | GPT-4.1 | OpenAI (2023a) | |
| 004 | GPT-4.1n | GPT-4.1-nano | OpenAI (2023b) | |
| 005 | GPT-o3-m | GPT-o3-mini | OpenAI (2024e) | |
| 006 | GPT-4o-m | GPT-4o-mini | OpenAI (2024c) | |
| 007 | Grok-3 | Grok-3 | xAI (2025) | |
| 008 | Claude-3.7S | Claude-3.7 Sonnet | Anthropic (2025) | |
| 009 | DS-chat | DeepSeek-Chat | AI (2024) | |
| Setting | Grader ID | Abbreviation | Full Name | Citation |
MultiTask |
001 | GPT-4o | GPT-4o | OpenAI (2024b) |
| 002 | GPT-4.1 | GPT-4.1 | OpenAI (2023a) | |
| 003 | GPT-o3 | GPT-o3 | OpenAI (2024d) | |
| 004 | GPT-o1 | GPT-o1 | OpenAI (2023c) | |
SingleTask |
001 | GPT-4o | GPT-4o | OpenAI (2024b) |
| 002 | GPT-4.1 | GPT-4.1 | OpenAI (2023a) | |
| 003 | GPT-o3 | GPT-o3 | OpenAI (2024d) | |
| 004 | GPT-o1 | GPT-o1 | OpenAI (2023c) |
@@ A.2 Formal Hypothesis Testing Framework for Wilcoxon Signed-Rank Significance Validity
To rigorously assess whether the Roundtable Policy (RT) provides statistically significant improvements over individual LLM baselines, we apply the non-parametric Wilcoxon signed-rank test to paired, round-level performance data. This test evaluates whether the median of the differences between two matched samples significantly exceeds zero, under the null hypothesis of no improvement.
@@
: score obtained by RT in round ,
@@
: relative gain of RT over player in round (as percentage).
@@
A.2.1 Application to MultiTask Evaluation
In the MultiTask evaluation, each task represents a distinct skill domain (e.g., math, geoscience, mechanics). Let be the set of benchmark tasks.
@@ A.2.1 Application to MultiTask Evaluation
We consider RT statistically superior to player if:
@@
A.2.2 Application to SingleTask Evaluation
In the SingleTask evaluation, each round produces rubric-based scores over 9 dimensions (covering three traits for each of Background, Methodology, and Impact).
@@ A.2.2 Application to SingleTask Evaluation
The MultiTask analysis evaluates whether RT consistently improves performance across structurally diverse problems and domains. The SingleTask analysis inspects RT’s advantage across nuanced dimensions of academic writing quality. In both cases, aggregate-level tests offer a robust indicator of RT’s overall statistical superiority over individual LLM baselines.
@@ A.3 Formal Framework for Kendall’s Tau Rank Correlation Analysis
Application to MultiTask.
In the MultiTask setting, each task yields a separate ranking of player models by each grader. For every round and task , we compute for all pairs of graders using their per-round rankings of the nine players. Averaging over all rounds gives
@@
Application to SingleTask.
In the SingleTask setting, each round produces rubric-based scores across nine dimensions . For each dimension , graders generate rankings of the players, and we compute for all pairs . Averaging across rounds yields
@@
In both MultiTask and SingleTask, Kendall’s Tau provides a principled measure of inter-grader consistency. The distinction lies in the evaluation unit: tasks for MultiTask versus rubric dimensions for SingleTask.
@@ Appendix B Detailed Demonstration of Roundtable Policy
To provide a concrete illustration of how Roundtable Policy (RT) operates in practice, we present two representative demonstrations: one in the MultiTask regime (Figure 7) and one in the SingleTask regime (Figure 6). These examples showcase how RT transforms fragmented outputs from multiple LLMs into coherent, factually grounded, and narratively rigorous responses.
MultiTask demonstration.
@@
Figure 7 presents a representative MultiTask round, where nine LLM players respond simultaneously to a heterogeneous set of scientific queries spanning mechanics, thermodynamics, semiconductor physics, mathematics, geoscience, and biology. Individual player outputs illustrate the limitations of single-model reasoning: some provide only partial explanations (e.g., noting that paper clips corrode without addressing stress relaxation), others introduce inaccuracies (e.g., oversimplifying nucleation–growth dynamics), and several fail to maintain consistency across domains (e.g., contradictory accounts of polymer glass transition). These fragmented outputs reflect the heterogeneity of LLM competencies—stronger players excel in certain areas, while weaker ones often produce vague or under-specified responses.
@@
By contrast, the fused response generated by RT demonstrates marked improvements in both factual coverage and reasoning depth. For instance, RT not only describes corrosion of paper clips but also contextualizes long-term degradation and mechanical implications; in the nucleation–growth problem, RT integrates complementary insights into thermodynamic driving forces and kinetic limitations, producing a textbook-quality explanation. Similarly, in the polymer glass transition question, RT combines molecular structure, chain length, and plasticizer effects into a unified account that no single player captured in full.
@@
Committee grading further substantiates these gains: RT exhibits higher factual accuracy and reduced uncertainty intervals across tasks. Improvements are particularly pronounced in symbolic and reasoning-intensive domains such as thermodynamics and mathematics, where RT delivers correct solutions while clarifying intermediate reasoning steps. These results highlight how RT balances correctness and clarity across heterogeneous domains.
SingleTask demonstration.
@@
Figure 6 illustrates the SingleTask regime, where each player generates a three-paragraph scientific proposal (Background, Methodology, Impact). Individual drafts vary widely in creativity, methodological depth, and logical flow. Some emphasize novelty but lack technical detail; others offer rigorous methods but fail to contextualize the problem or articulate downstream impact. Such stylistic and structural inconsistencies exemplify the difficulty of producing a coherent scientific narrative with a single LLM.
@@
In contrast, the fused proposal produced by RT exhibits a markedly stronger narrative arc. The Background situates the research problem within its broader scientific context; the Methodology clearly articulates technical contributions with precision; and the Impact section highlights both scientific significance and societal relevance. Grading results confirm consistent improvements across all rubric dimensions—creativity, scientific rigor, and logical coherence—demonstrating that RT effectively mitigates individual stylistic biases while amplifying collective strengths. Even strong baseline players benefit: while their drafts are often polished in one dimension, RT integrates complementary insights to deliver uniformly high-quality writing.
@@
Together, these demonstrations show how Roundtable Policy transforms fragmented or uneven model outputs into coherent, high-quality responses. In the MultiTask regime, RT enhances cross-domain reasoning reliability by integrating complementary knowledge across scientific fields. In the SingleTask regime, RT substantially improves narrative structure and explanatory depth in long-form proposal writing. These examples reinforce the central claim that RT enables LLM ensembles to achieve more factual, rigorous, and narratively coherent outputs in scientific applications.

### v2

@@ 4 Related Work
Multi-Agent Systems. Early approaches show that majority (Wang et al., 2023) and confidence-weighted schemes (Taubenfeld et al., 2025) can improve reliability by combining diverse reasoning trajectories. Besides, Du et al. (2024) present multi-agent debate, while Liang et al. (2024) introduce asymmetric debater–judge roles in debate. Further collaborative paradigm emphasize fine-grained interaction protocols (Chan et al., 2024; Zhang et al., 2024; Wang et al., 2025). Our work provides a complementary perspective by studying how reliability-aware consensus can be used during inference, compatible with other collaboration strategies.
LLMs for Science. Large foundation models have driven broad scientific advances, from protein structure prediction (Senior et al., 2020; Jumper et al., 2021; Baek et al., 2021; Lin et al., 2023; Elnaggar et al., 2021) to domain-specialized tasks spanning math, chemistry, biology, and geoscience (Romera-Paredes et al., 2024; Trinh et al., 2024; Yue et al., 2024; Yu et al., 2024; Labrak et al., 2024; Li et al., 2023). Our work complements increasingly complex agentic workflows that couple LLMs with external tools and experimental pipelines (Yao et al., 2023; Boiko et al., 2023; Lu et al., 2024; Yamada et al., 2025) by focusing on end-to-end inference-time reasoning and aggregation.
Benchmarks for Scientific Reasoning. Scientific reasoning has been widely benchmarked. Math GSM8K (Cobbe et al., 2021) focus on step-by-step numerical reasoning, biomedical PubMedQA (Jin et al., 2019) assess Q&A grounded in literature, GPQA and AlphaGeometry (Trinh et al., 2024) introduce problem-solving in specialized settings. Multi-domain benchmarks such as MMLU (Hendrycks et al., 2021) and BIG-Bench (bench authors, 2023) evaluate general reasoning across a broad range of technical subjects, ARC (Clark et al., 2018), GSM-Symbolic (Mirzadeh et al., 2025), and GPQA-Diamond (Rein et al., 2024) further emphasize compositional reasoning in diverse settings. Our work extends these efforts by integrating heterogeneous, long-context scientific reasoning, providing explicit dimensions for assessing expertise in multi-agent systems.
5 Conclusion
We presented Roundtable Policy, an inference-time reasoning framework for multi-agent systems, and demonstrated its efficacy through extensive experiments. More broadly, this work suggests that multi-agent collaboration for complex tasks may benefit from explicitly modeling and stabilizing agent reliability and consensus, motivating principled inference-time strategies, and evaluation frameworks that better reflect the characteristics of downstream tasks.
Impact Statement
Roundtable Policy can boost performance in diverse downstream tasks using multi-agent systems.
@@ A.1 Symbols and Abbreviations
: the number of player agents (proposer agents, pretrained LLMs generating candidate responses). In our experiments .
@@
: the fusion agent (aggregator agent), which synthesizes a final consensus answer from all player responses using the confidence-weight table.
@@
: the corresponding reference answers (available only in ScienceEval).
@@
: the fused consensus response in round , using a pre-constructed confidence-weight table.
| Setting | Player ID | Abbreviation | Full Name |
| SciEval | 001 | GPT-4o | GPT-4o |
| 002 | Grok-2 | Grok-2 | |
| 003 | Claude-3.5H | Claude-3.5 Haiku | |
| 004 | GPT-4.5p | GPT-4.5-preview | |
| 005 | DS-chat | DeepSeek-Chat | |
| 006 | Claude-3.7S | Claude-3.7 Sonnet | |
| 007 | Llama-3.3 | Llama-3.3 | |
| 008 | Qwen+ | Qwen-plus | |
| 009 | Gemini-2.0 | Gemini-2.0 | |
| SciNarrative | 001 | GPT-4o | GPT-4o |
| 002 | GPT-o3 | GPT-o3 | |
| 003 | GPT-4.1 | GPT-4.1 | |
| 004 | GPT-4.1n | GPT-4.1-nano | |
| 005 | GPT-o3-m | GPT-o3-mini | |
| 006 | GPT-4o-m | GPT-4o-mini | |
| 007 | Grok-3 | Grok-3 | |
| 008 | Claude-3.7S | Claude-3.7 Sonnet | |
| 009 | DS-chat | DeepSeek-Chat | |
| Setting | Grader ID | Abbreviation | Full Name |
| SciEval | 001 | GPT-4o | GPT-4o |
| 002 | GPT-4.1 | GPT-4.1 | |
| 003 | GPT-o3 | GPT-o3 | |
| 004 | GPT-o1 | GPT-o1 | |
| SciNarrative | 001 | GPT-4o | GPT-4o |
| 002 | GPT-4.1 | GPT-4.1 | |
| 003 | GPT-o3 | GPT-o3 | |
| 004 | GPT-o1 | GPT-o1 | |
| Setting | Aggregator Type | Full Name | |
| SciEval | Strong | GPT-4o | |
| Intermediate-1 | Llama-3.3 | ||
| Intermediate-2 | Qwen-Plus | ||
| Weak | Gemini-2.0 | ||
| SciNarrative | Strong | Claude-3.7S | |
| Intermediate-1 | Grok-3 | ||
| Intermediate-2 | GPT-4.1 | ||
| Weak | GPT-4o | ||
@@ A.2 Formal Hypothesis Testing Framework for Wilcoxon Signed-Rank Significance Validity
To rigorously assess whether the Roundtable Policy (RP) provides statistically significant improvements over individual LLM baselines, we apply the non-parametric Wilcoxon signed-rank test to paired, round-level performance data. This test evaluates whether the median of the differences between two matched samples significantly exceeds zero, under the null hypothesis of no improvement.
@@
: score obtained by RP in round ,
@@
: relative gain of RP over player in round (as percentage).
@@
A.2.1 Application to ScienceEval Evaluation
In the ScienceEval evaluation, each task represents a distinct skill domain (e.g., math, geoscience, biology). Let be the set of benchmark tasks.
@@ A.2.1 Application to ScienceEval Evaluation
We consider RP statistically superior to player if:
@@
A.2.2 Application to ScienceNarrativeEvaluation
In the ScienceNarrative evaluation, each round produces rubric-based scores over 9 dimensions (covering three traits for each of Background, Methodology, and Impact).
@@ A.2.2 Application to ScienceNarrativeEvaluation
The ScienceEval analysis evaluates whether RP consistently improves performance across structurally diverse problems and domains. The ScienceNarrative analysis inspects RP’s advantage across nuanced dimensions of academic writing quality. In both cases, aggregate-level tests offer a robust indicator of RP’s overall statistical superiority over individual LLM baselines.
@@ A.3 Formal Framework for Kendall’s Tau Rank Correlation Analysis
Application to ScienceEval.
In the ScienceEval setting, each task yields a separate ranking of player models by each grader. For every round and task , we compute for all pairs of graders using their per-round rankings of the nine players. Averaging over all rounds gives
@@
Application to ScienceNarrative.
In the ScienceNarrative setting, each round produces rubric-based scores across nine dimensions . For each dimension , graders generate rankings of the players, and we compute for all pairs . Averaging across rounds yields
@@
In both settings, Kendall’s Tau provides a principled measure of inter-grader consistency. The distinction lies in the evaluation unit: tasks for ScienceEval versus rubric dimensions for ScienceNarrative.
@@ Appendix B Detailed Demonstration of Roundtable Policy
To provide a concrete illustration of how Roundtable Policy (RP) operates in practice, we present two representative demonstrations: one in the ScienceEval regime (Figure 9) and one in the ScienceNarrative regime (Figure 8). These examples showcase how RP transforms fragmented outputs from multiple LLMs into coherent, factually grounded, and narratively rigorous responses.
ScienceEval demonstration.
@@
Figure 9 presents a representative ScienceEval round, where nine LLM players respond simultaneously to a heterogeneous set of scientific queries spanning mechanics, thermodynamics, semiconductor physics, mathematics, geoscience, and biology. Individual player outputs illustrate the limitations of single-model reasoning: some provide only partial explanations (e.g., noting that paper clips corrode without addressing stress relaxation), others introduce inaccuracies (e.g., oversimplifying nucleation–growth dynamics), and several fail to maintain consistency across domains (e.g., contradictory accounts of polymer glass transition). These fragmented outputs reflect the heterogeneity of LLM competencies—stronger players excel in certain areas, while weaker ones often produce vague or under-specified responses.
@@
By contrast, the fused response generated by RP demonstrates marked improvements in both factual coverage and reasoning depth. For instance, RP not only describes corrosion of paper clips but also contextualizes long-term degradation and mechanical implications; in the nucleation–growth problem, RP integrates complementary insights into thermodynamic driving forces and kinetic limitations, producing a textbook-quality explanation. Similarly, in the polymer glass transition question, RP combines molecular structure, chain length, and plasticizer effects into a unified account that no single player captured in full.
@@
Committee grading further substantiates these gains: RP exhibits higher factual accuracy and reduced uncertainty intervals across tasks. Improvements are particularly pronounced in symbolic and reasoning-intensive domains such as thermodynamics and mathematics, where RP delivers correct solutions while clarifying intermediate reasoning steps. These results highlight how RP balances correctness and clarity across heterogeneous domains.
ScienceNarrative demonstration.
@@
Figure 8 illustrates the ScienceNarrative regime, where each player generates a three-paragraph scientific proposal (Background, Methodology, Impact). Individual drafts vary widely in creativity, methodological depth, and logical flow. Some emphasize novelty but lack technical detail; others offer rigorous methods but fail to contextualize the problem or articulate downstream impact. Such stylistic and structural inconsistencies exemplify the difficulty of producing a coherent scientific narrative with a single LLM.
@@
In contrast, the fused proposal produced by RP exhibits a markedly stronger narrative arc. The Background situates the research problem within its broader scientific context; the Methodology clearly articulates technical contributions with precision; and the Impact section highlights both scientific significance and societal relevance. Grading results confirm consistent improvements across all rubric dimensions—creativity, scientific rigor, and logical coherence—demonstrating that RP effectively mitigates individual stylistic biases while amplifying collective strengths. Even strong baseline players benefit: while their drafts are often polished in one dimension, RP integrates complementary insights to deliver uniformly high-quality writing.
@@
Together, these demonstrations show how Roundtable Policy transforms fragmented or uneven model outputs into coherent, high-quality responses. In the ScienceEval regime, RP enhances cross-domain reasoning reliability by integrating complementary knowledge across scientific fields. In the ScienceNarrative regime, RP substantially improves narrative structure and explanatory depth in long-form proposal writing. These examples reinforce the central claim that RP enables LLM ensembles to achieve more factual, rigorous, and narratively coherent outputs in scientific applications.